200 mL of water sample has hardness equivalent to 25 mL of 0.08 N MgSO4. What is the hardness of the sample?
1
Expert's answer
2020-06-25T07:44:51-0400

200 ml of water and density of water is 1 gm/ml .

Mass of water is 200 gram .


number of gram equivalent of MgSO4 = number of gram equivalent of CaCO3 .


So , number of gram equivalent of CaCO3 = 0.002 gm equivalent...

moles of CaCO3 =( 0.002 / 2 ) moles .

mass of CaCO3 = 0.1 gm...

where mass = moles * molar mass ( molar mass of CaCO3 = 100 gm/mol ).


Degree of Hardness = (Mass of CaCO3 ) * 106 / Mass of water .

Degree of Hardness = 0.1 *106 / 200 = 500 ppm . answer.